The sum of two numbers is 22. Using q as the smaller number, translate "five less than the larger number" into a variable expression and simplify.
Answer by stanbon(75887)

The sum of two numbers is 22. Using q as the smaller number, translate "five less than the larger number" into a variable expression and simplify.
Equations:
smaller: q
larger: p
p + q = 22
q = p-5
-----
Substitute for "q" and solve for "p":
p + p-5 = 22
2p = 27
p = 13.5
q = p-5 = 8.5
===================
Cheers,
Stan H.